If 6.68 g of CuNO3 is dissolved in water to make a 0.520 M solution, what is the volume of the solution?
1
Expert's answer
2019-04-15T07:54:22-0400

"n(Cu(NO3)2)=\dfrac{6.68g}{187.57g/mol}=0.0356moles"

"V=\dfrac{n}{c}=\dfrac{0.0356moles}{0.52mol/L}=0.06846L=68.46mL"